A little more information about the IMC loop tweak...
I increased the overall IMC loop gain by 4dB, and decreased the FAST gain (which determines the PZT/EOM crossover) by 3dB. This changed the AO transfer function from the blue trace to the green trace in the first plot. This changed the CARM loop open loop TF shape from the unfortunate blue shape to the more pleasing green shape in the second plot. The red trace is the addition of one super boost.
 
Oddly, these transfer functions look a bit different than what I measured in March (ELOG 11167), which itself differed from the shaping done December of 2014 (ELOG 10841).
I haven't yet attempted any 1F handoff of the PRMI since relocking, but back when Jenne and I did so in April, the lock was definitely less stable. My suspicion is that we may need more CARM supression; we never computed the loop gain requirement that ensures that the residual CARM fluctuations witnessed by, say, REFL55 are small enough to use as a reliable PRMI sensor.
I should be able to come up with this with data from last night. |